OpenStudy (anonymous):

Input the equation of the given line in standard form. The line through (1, -4) and parallel to 2x + 3y = 4.

OpenStudy (texaschic101):

do you know how to find the slope in the equation 2x + 3y = 4 ?

OpenStudy (texaschic101):

put the equation in y = mx + b form, and m will be your slope. Can you find the slope ?

OpenStudy (anonymous):

would the slope me 2/3?

OpenStudy (anonymous):

be

OpenStudy (texaschic101):

ok....2x + 3y = 4 3y = -2x + 4 y = -2/3x + 4/3 the slope (the m position) is -2/3 A parallel line will have the same slope. Now we will use y = mx + b slope(m) = -2/3 (1,-4) x = 1 and y = -4 now we sub and solve for b -4 = -2/3(1) + b -4 = 2/3 + b -4 - 2/3 = b -12/3 - 2/3 = b -14/3 = b your parallel equation is : y = -2/3x - 14/3, but we need it in standard form Ax + By = C y = -2/3x - 14/3 -- multiply by 3 3y = -2x - 14 -- add 2x to both sides 2x + 3y = -14 ====> standard form any questions ?

OpenStudy (anonymous):

No. Thank you so much! :)

OpenStudy (texaschic101):

wait...I think I messed up...hold on

OpenStudy (texaschic101):

-4 = -2/3(1) + b -4 = -2/3 + b -4 + 2/3 = b -12/3 + 2/3 = b -10/3 = b y = -2/3x - 10/3 -- multiply by 3 3y = -2x - 10 -- add 2x 2x + 3y = -10 --->standard form sorry about that

OpenStudy (texaschic101):

thats better :)